OpenStudy (abbycross167):

Can someone please help me with this math question? I will give a medal and will fan... Question - Simplify (-3a^2 + 8a – 2) – (-4a^2 – 2a + 6)

OpenStudy (xguardians):

1. Distribute the negative. 2. Combine "like" terms.

OpenStudy (abbycross167):

What do you mean Distribute the negative?

OpenStudy (superpanda1022):

Distribute the negative means distribute -1(-4a^2-2a+6)

OpenStudy (superpanda1022):

Then make them add each other

OpenStudy (superpanda1022):

-3a^2 + 8a – 2+4a^2+2a-6 Combine like terms

OpenStudy (abbycross167):

so -3a^2+4a^2 =7a^2 8a+2a = 10a 2+6=8

OpenStudy (superpanda1022):

-3a^2+4a^2=a^2 8a+2a=10a -2-6=-8 If you don't know how to add negatives just use calculator

OpenStudy (abbycross167):

oh ok... I did use a calculator but I see my mistake

OpenStudy (superpanda1022):

Don't forget the negatives next time! :)

OpenStudy (abbycross167):

yes sir/ma'am :) thank you for helping me

OpenStudy (superpanda1022):

Not a problem!